If B is an angle in standard position and (3,5) are the coordiantes of a point on the terminal are mof B, then cos B is?
Answer by eperette(173)

cosB=x/r
x^2 + y^2 =r^2
(3)^2 + (5)^2 = r^2
9 + 25 = r^2
34= r^2
r= +- (sqr (34))
But remember r is always positive...therefore r= sqr(34)
SO
cosB=3/sqr(34)
Remember it is not mathematically accepted to have a square root in denominator, so lets get rid of it:
3/sqr(34) = (3/sqr(34))* (sqr(34)/sqr(34))= 3sqr(34)/34
*****
cosB= 3sqr(34)/34